OpenStudy (anonymous):

7x2 + 10x = -3 A. 3/7, -1 B. -3/7 C. -3/7, 1 D. -3/7, -1 E. None of the above

OpenStudy (saylilbaby):

D

OpenStudy (kamibug):

Add three to both sides to set the equation equal to zero. Then factor 7x^2 + 10x +3. Here's a good factoring calculator: http://www.freemathhelp.com/factoring-calculator.php Once you have the factors, set each one equal to zero and solve for the variable. :)

OpenStudy (jordanloveangel):

@$Young_Money$ 7x2-10x-8=0 7x2 -10x -8 = 0 ax2 bx c = 0 a = 7 b = -10 c = -8 Δ = b2-4ac Δ = -10² - (4 * 7 * -8) Δ = 100 - (4*-56) Δ = 100 - (-224) Δ = 324 Δ > 0 7x2-10x-8=0 x = -b ∓ √Δ 2a x' = 10 - √324 2*7 = 10 - 18.0 14 = -8.0 14 = -0.571428571429 x" = 10 + √324 2*7 = 10 + 18.0 14 = 28.0 14 = 2.0 x' = -28.0 x" = 98.0 S = {x ∈ R / x' = -28.0, x" = 98.0 }